# Heegaard Floer Surgery Formula and Cosmetic Surgeries

### v1

@@ 5. Cosmetic Surgeries
Let $K\subset Y$ be a blow-up unknotted knot, let $p,q>0$ coprime. Let $b=\operatorname{rank}(\widehat{B})=\operatorname{rank}\widehat{HF}(Y)$. Then
(5.9) $\begin{split}\operatorname{rank}\widehat{HF}(Y_{p/q}(K))&=q(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &+2q\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})+2t_{K}^{p/q}-pb,\end{split}$
@@
(5.10) $t_{K}^{p/q}$ $=\sum_{j=0}^{p-1}\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right).$
@@
By Theorem 4.1, $\widehat{HF}(Y_{p/q}(K))\cong H_{*}(\widehat{\mathbb{X}}_{p/q})$, the homology of the mapping cone of $\widehat{D}_{p/q}$ from $\widehat{{\mathbb{A}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{A}}_{i}$ to $\widehat{{\mathbb{B}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{B}}_{i}$. We abuse notation to let $\widehat{A}_{j/q}$ denote $\widehat{A}_{\lfloor j/q\rfloor}$, and also $\widehat{B}_{j/q}$, $\widehat{v}_{j/q}$, $\widehat{h}_{j/q}$ similarly. So
$\widehat{\mathbb{A}}=\bigoplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\prod_{r\in{\mathbb{Z}}}\widehat{A}_{(i+pr)/q}=\prod_{j\in{\mathbb{Z}}}\widehat{A}_{j/q}$
has $q$ copies of $\widehat{A}_{s}$ for each $s\in{\mathbb{Z}}$, and under the map $\widehat{v}$, each $\widehat{A}_{j/q}$ corresponds to a copy of $\widehat{B}$ in $\widehat{\mathbb{B}}$, which we shall denote $\widehat{B}_{j/q}$. Then $\widehat{h}_{j/q}$ maps $\widehat{A}_{j/q}$ to $\widehat{B}_{(j+p)/q}$. Thus, for each $s\in{\mathbb{Z}}$, there are $p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s-1}$ and $q-p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s}$.
@@
and $\widehat{D}_{p/q}^{c}$ is the restriction of $\widehat{D}_{p/q}$ to $\widehat{A}_{p/q}^{c}$, with the exception that for any $x\in\widehat{A}_{j/q}$ with $-qc+1\leq j\leq-qc+p$, $\widehat{D}_{p/q}^{c}(x)=\widehat{h}_{j/q}(x)$, and for any $x\in\widehat{A}_{j/q}$ with $qc-p\leq j\leq qc-1$, $\widehat{D}_{p/q}^{c}(x)=\widehat{v}_{j/q}(x)$. Note that it suffices to take $c\geq g+p/q+1$, and
@@
We construct a basis for $*\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}$. For $j\geq 0$, let
$*\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$
be a basis of $*\ker(\widehat{v}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose an $\alpha_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\alpha_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$. For $j<0$, let
$*\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{h}_{j/q})_{*}}$
be a basis of $*\ker(\widehat{h}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose a $\beta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\beta_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$. Lastly, for $10\leq j\leq p-$ 1, let
@@
be a basis of the subspace of $H_{*}(\widehat{A}_{j/q})$ that is mapped isomorphically to $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$ under $(\widehat{v}_{j/q})_{*}$. Choose $z_{j/q}^{i}\in H_{*}(\widehat{A}_{(j-p)/q})$ such that $(\widehat{v}_{j/q})_{*}(y_{j/q}^{i})=(\widehat{h}_{(j-p)/q})_{*}(z_{j/q}^{i})$. Choose $\gamma_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ and $\delta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{y}_{j/q}^{i}=y_{j/q}^{i}+z_{j/q}^{i}+\gamma_{j/q}^{i}+\delta_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$.
@@
is linearly independent, and it is finite by Lemma 5.3. Moreover, it spans $*\ker(\widehat{D}_{p/q})_{*}$: let $*z\in\ker(\widehat{D}_{p/q})_{*}$ be nonzero, and suppose first that $z\in H_{*}(\oplus_{k\geq 0}\widehat{A}_{k/q})$. Let $z_{j/q}$ denote the term of $z$ in $H_{*}(\widehat{A}_{j/q})$, and let
@@
Then $z\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, so $*z_{j/q}\in\operatorname{Span}\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$: otherwise if $*z_{j/q}\notin\ker(\widehat{v}_{j/q})_{*}$, then $*z\notin\ker(\widehat{D}_{p/q})_{*}$ since $z\in H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$. Thus,
@@
[…]
@@
[…]
@@
[…]
Now for the general case $*z\in\ker(\widehat{D}_{p/q})_{*}$, we claim that for each $10\leq j\leq p-$ 1,
@@
Indeed, suppose this is not true for some $10\leq j\leq p-$ 1, then $*(\widehat{v}_{j/q})_{*}(z_{j/q})\notin\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, so $(\widehat{D}_{p/q})_{*}(z)\neq 0$, a contradiction. Thus,
@@
Therefore, $S$ is a basis for $*\ker(\widehat{D}_{p/q})_{*}$, so since $*\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}\ker(\widehat{h}_{-s})_{*}$,
@@
Since $*\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})-\operatorname{rank}\ker(\widehat{v}_{s})_{*}$, $\operatorname{rank}H_{*}(\widehat{A}_{s})=b$ for all $|s|\geq g$, and $*\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}=\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})-\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})+\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*}$ for some large $c$, it follows that
$*\displaystyle\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$ $=b(2c-1)q-pb-q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s})$
@@
and therefore Equation 5.9 holds.
@@
Since $H_{1}(Y_{\pm p/q}(K))\cong H_{1}(Y)\oplus{\mathbb{Z}}/p{\mathbb{Z}},$ we can fix $p>0$ throughout. First, let $q$ be relatively prime to $p$ and $q\geq p$. Then $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}=\operatorname{Im}(\widehat{v}_{0})_{*}$ for all $j=0,\dots,p-1$. Thus $*t_{K}^{p/q}=p\operatorname{rank}(\widehat{v}_{0})_{*}$. Therefore, according to equation 5.9, for positive integral $q$ with $0<p/q\leq 1$, $\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ is a nondecreasing function of $q$. The function is strictly increasing unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Hence, by Proposition 5.5, $K$ is the unknot.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result is true whenever $0<r<s\leq 1$.
Now suppose $q,q^{\prime}$ are relatively prime to $p$ and $0<q<p<q^{\prime}$. Then
$t_{K}^{p/q}-t_{K}^{p/q^{\prime}}$ $*\displaystyle\leq\sum_{j=0}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-p\operatorname{rank}(\widehat{v}_{0})_{*}$
$*\displaystyle=\sum_{j=q}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-(p-q)\operatorname{rank}(\widehat{v}_{0})_{*}$
$\leq(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*}).$
Therefore,
@@
$+2t_{K}^{p/q^{\prime}}-2t_{K}^{p/q}$
$\geq 2(q^{\prime}-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})-2(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})$
@@
Let $K\subset Y$ be a nontrivial blow-up unknotted knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $g(K)\geq 2$ or $\operatorname{rank}H_{*}(\widehat{A}_{0})>b$, then $Y_{1}(K)\ncong\pm Y$.
@@
$\operatorname{rank}\widehat{HF}(Y_{1}(K))\geq\operatorname{rank}(\widehat{v}_{0})_{*}+\operatorname{rank}\ker(\widehat{v}_{0})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{0})\geq b,$
and it is strictly greater unless $\operatorname{rank}H_{*}(\widehat{A}_{0})=b$ and $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 1$, which implies $g(K)\leq 1$.

### v2

@@ 5 Cosmetic Surgeries
Let $K\subset Y$ be a blow-up bounding knot, let $p,q>0$ coprime. Let $b=\operatorname{rank}(H_{*}(\widehat{B}))=\operatorname{rank}\widehat{HF}(Y)$. Then
$\begin{split}\operatorname{rank}\widehat{HF}(Y_{p/q}(K))&=q(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &+2q\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})+2t_{K}^{p/q}-pb,\end{split}$ (5.9)
@@
$t_{K}^{p/q}$ $=\sum_{j=0}^{p-1}\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right).$ (5.10)
@@
By Theorem 4.1, $\widehat{HF}(Y_{p/q}(K))\cong H_{*}(\widehat{\mathbb{X}}_{p/q})$, the homology of the mapping cone of $\widehat{D}_{p/q}$ from $\widehat{{\mathbb{A}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{A}}_{i}$ to $\widehat{{\mathbb{B}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{B}}_{i}$ (we have omitted the slope $p/q$ from the notation and summed over all $\operatorname{Spin}^{c}$ structures on $Y$). So
$\widehat{\mathbb{A}}=\bigoplus_{i=0}^{p-1}\prod_{r\in{\mathbb{Z}}}\widehat{A}_{(i+pr)/q}=\prod_{j\in{\mathbb{Z}}}\widehat{A}_{j/q}$
has $q$ copies of $\widehat{A}_{s}$ for each $s\in{\mathbb{Z}}$, and under the map $\widehat{v}$, each $\widehat{A}_{j/q}$ corresponds to a copy of $\widehat{B}$ in $\widehat{\mathbb{B}}$, which we shall denote $\widehat{B}_{j/q}$. Then $\widehat{h}_{j/q}$ maps $\widehat{A}_{j/q}$ to $\widehat{B}_{(j+p)/q}$. Thus, for each $s\in{\mathbb{Z}}$, there are $p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s-1}$ and qpq-p copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s}$.
@@
and $\widehat{D}_{p/q}^{c}$ is the restriction of $\widehat{D}_{p/q}$ to $\widehat{A}_{p/q}^{c}$, with the exception that for any $x\in\widehat{A}_{j/q}$ with $qc+1\leq j\leq-qc+p$, $\widehat{D}_{p/q}^{c}(x)=\widehat{h}_{j/q}(x)$, and for any $x\in\widehat{A}_{j/q}$ with $qc-p\leq j\leq qc-1$, $\widehat{D}_{p/q}^{c}(x)=\widehat{v}_{j/q}(x)$. Note that it suffices to take $c\geq g+p/q+1$, and
@@
We construct a basis for $\ker(\widehat{D}_{p/q})_{*}$ of the entire mapping cone, which has the same rank as $\ker(\widehat{D}_{p/q}^{c})_{*}$ for large $c$ by Lemma 2.6. For $j\geq 0$, let
$\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$
be a basis of $\ker(\widehat{v}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose an $\alpha_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\alpha_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$. For $j<0$, let
$\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{h}_{j/q})_{*}}$
be a basis of $\ker(\widehat{h}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose a $\beta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\beta_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$. Lastly, for $0\leq j\leq p-1$, let
@@
be a basis of a subspace of $H_{*}(\widehat{A}_{j/q})$ that is mapped isomorphically to $\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$ under $(\widehat{v}_{j/q})_{*}$. Choose $z_{j/q}^{i}\in H_{*}(\widehat{A}_{(j-p)/q})$ such that $(\widehat{v}_{j/q})_{*}(y_{j/q}^{i})=(\widehat{h}_{(j-p)/q})_{*}(z_{j/q}^{i})$. Choose $\gamma_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ and $\delta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{y}_{j/q}^{i}=y_{j/q}^{i}+z_{j/q}^{i}+\gamma_{j/q}^{i}+\delta_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$.
@@
is linearly independent, and it is finite by Lemma 2.6. Moreover, it spans $\ker(\widehat{D}_{p/q})_{*}$: let $z\in\ker(\widehat{D}_{p/q})_{*}$ be nonzero, and suppose first that $z\in H_{*}(\oplus_{k\geq 0}\widehat{A}_{k/q})$. Let $z_{j/q}$ denote the term of $z$ in $H_{*}(\widehat{A}_{j/q})$, and let
@@
Then $z\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, so $z_{j/q}\in\operatorname{Span}\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$: otherwise if $z_{j/q}\notin\ker(\widehat{v}_{j/q})_{*}$, then since $z\in H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, we have $z_{(j-p)/q}=0$ so
$(\widehat{v}_{j/q})_{*}(z_{j/q})+(\widehat{h}_{(j-p)/q})_{*}(z_{(j-p)/q})\neq 0\in H_{*}(\widehat{B}_{j/q})$
which means $z\notin\ker(\widehat{D}_{p/q})_{*}$. Thus, we can subtract some combination of the basis elements $\widetilde{x}_{j/q}^{i}$ to eliminate the $H_{*}(\widehat{A}_{j/q})$ component of $z$:
@@
for some indexing set $J\subset\{1,\dots,\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}\}$. Inductively, we can then write
@@
for some finite indexing sets $I\subset\{k\geq j\}$, $J_{k}\subset\{1,\dots,\operatorname{rank}\ker(\widehat{v}_{k/q})_{*}\}$. Similarly, if $z\in H_{*}(\oplus_{k<0}\widehat{A}_{k/q})$, then we can write
@@
for some finite indexing set $I\subset\{k\leq j\}$, $J_{k}\subset\{1,\dots,\operatorname{rank}\ker(\widehat{h}_{k/q})_{*}\}$.
Now for the general case $z\in\ker(\widehat{D}_{p/q})_{*}$, we claim that for each $0\leq j\leq p-1$,
@@
Indeed, suppose this is not true for some $0\leq j\leq p-1$, then $(\widehat{v}_{j/q})_{*}(z_{j/q})\notin\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, so $(\widehat{D}_{p/q})_{*}(z)\neq 0$, a contradiction. Thus,
@@
Therefore, $S$ is a basis for $\ker(\widehat{D}_{p/q})_{*}$, so since $\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}\ker(\widehat{h}_{-s})_{*}$ by the hat version of Lemma 2.5 and these are 0 when $s\geq g$ by Lemma 2.6,
@@
Since $\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})-\operatorname{rank}\ker(\widehat{v}_{s})_{*}$ and $\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}=\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})-\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})+\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*}$ for some large $c$, it follows that
$\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$ $=b(2c-1)q-pb-q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s})$
@@
where the second equality uses that $\operatorname{rank}H_{*}(\widehat{A}_{s})=b$ for all $|s|\geq g$ by Lemmas 2.5 and  2.6. Therefore, because
$\operatorname{rank}(H_{*}(\widehat{\mathbb{X}}_{p/q}))=\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}+\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q})_{*}$
and
$\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}=\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*},\hskip 14.22636pt\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q})_{*}=\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$
for large enough $c$, equation 5.9 holds.
@@
Let $K\subset Y$ be a blow-up bounding knot, and suppose $0<r<s\leq 1$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$. Then $K$ is the unknot. Furthermore, if $K\subset Y$ is a blow-up unknotted knot and $0<r<s$ are distinct positive rational numbers such that $Y_{r}(K)\cong\pm Y_{s}(K)$, then $r,s>1$ or $K$ is the unknot.
@@
Let $K\subset Y$ be blow-up bounding. Since $H_{1}(Y_{\pm p/q}(K))\cong H_{1}(Y)\oplus{\mathbb{Z}}/p{\mathbb{Z}},$ we can fix $p>0$ throughout. First, let $q$ be relatively prime to $p$ and $q\geq p$. Then $\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}=\operatorname{Im}(\widehat{v}_{0})_{*}$ for all $j=0,\dots,p-1$. Thus $t_{K}^{p/q}=p\operatorname{rank}(\widehat{v}_{0})_{*}$. Therefore, according to equation 5.9, for positive integral $q$ with $0<p/q\leq 1$, $\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ is a nondecreasing function of $q$. The function is strictly increasing unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Hence, by Proposition 5.5, $K$ is the unknot.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the first part of the theorem is true. Also, the second part is true for $0<r<s\leq 1$ since blow-up unknotted knots are blow-up bounding.
Now suppose $K$ is blow-up unknotted and $q,q^{\prime}$ are relatively prime to $p$ and $0<q<p<q^{\prime}$. Then, with $)b=\operatorname{rank}(H_{*}(\widehat{B}$)),
$\begin{split}t_{K}^{p/q}-t_{K}^{p/q^{\prime}}&\leq\sum_{j=0}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-p\operatorname{rank}(\widehat{v}_{0})_{*}\\ &=\sum_{j=q}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-(p-q)\operatorname{rank}(\widehat{v}_{0})_{*}\\ &\leq(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &=0\end{split}$ (5.12)
since $(\widehat{v}_{0})_{*}$ is surjective by Lemma 5.4. By equation 5.9,
@@
$+2t_{K}^{p/q^{\prime}}-2t_{K}^{p/q}.$
The entire sum from $s=1$ to g1g-1 on the right hand side is nonnegative, so we ignore this term, and by Lemma 5.4, we have
$\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*}=\operatorname{rank}\ker(\widehat{v}_{0})_{*}.$
Together with inequality 5.12, we get
$\operatorname{rank}\widehat{HF}(Y_{p/q^{\prime}}(K))-\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ $\geq(q^{\prime}-q)\operatorname{rank}\ker(\widehat{v}_{0})_{*}$
@@
Let $K\subset Y$ be a nontrivial blow-up bounding knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $K$ is a nontrivial blow-up unknotted knot, then $Y_{1}(K)\ncong\pm Y$.
@@
$\operatorname{rank}\widehat{HF}(Y_{1}(K))\geq\operatorname{rank}(\widehat{v}_{0})_{*}+\operatorname{rank}\ker(\widehat{v}_{0})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{0})\geq b$
by Lemma 5.4 since $K$ is blow-up unknotted, and it is strictly greater unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$, which implies $K$ is the unknot.
@@
Acknowledgements The author would like to thank Yi Ni for advising him in this project and the referee for many helpful comments.